OpenStudy (anonymous):

Please Help! Could someone tell me what the answer is, to this question? Thanks! "Multiply: (x-4)/(2x-18) x (x^2-7x-18)/(x^2-16)"

OpenStudy (campbell_st):

do some factorising \[\frac{x -4}{2(x-9)} \times \frac{(x - 9)(x +2)}{(x+4)(x-4)}\] eliminating common factors leaves (x -9) and (x - 4) \[\frac{(x +2)}{2(x+4)}\]

OpenStudy (anonymous):

Aha! Just what I thought! Thanks so much , campbell_st! I just wanted to check, to see if my math curriculum was incorrect - I tried this problem twice, and I got the same answer you did.